# Lightweight and Direct Document Relevance Optimization for Generative Information Retrieval

### v1

@@ Abstract.
submissionid: 821copyright: rightsretainedjournalyear: 2025conference: The 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13-18, 2025; Padua, Italyisbn: ccs: Information systems Retrieval models and rankingccs: Information systems Learning to rankccs: Information systems Language models
@@ 1. Introduction
Building on the success of transformer-based pre-trained language models, recent research has explored various neural retrieval approaches (Yates et al., 2021): learned sparse retrieval (Formal et al., 2021a, b), dense retrieval (Karpukhin et al., 2020a; Xiong et al., 2020; Zeng et al., 2022; Zhan et al., 2021), and cross-encoders (Nogueira and Cho, 2019). A new paradigm has recently been added to this palette, generative information retrieval (GenIR(Tay et al., 2022; Metzler et al., 2021). This approach uses pre-trained encoder-decoder models as differentiable search indexes (DSI). It has inspired the development of several models, including NCI (Wang et al., 2022), DSI-QG (Zhuang et al., 2022), DSI++ (Mehta et al., 2023), SEAL (Bevilacqua et al., 2022), ULTRON (Zhou et al., 2022b), TOME (Ren et al., 2023), IncDSI (Kishore et al., 2023), RIPOR (Zeng et al., 2024b), and PAG (Zeng et al., 2024a).
@@
GenIR models represent documents as sequences of unique document identifiers (docids), generated autoregressively, where each token is conditioned on the query encoding and previously generated tokens. The generation process is controlled through (constrained) beam search (Mehta et al., 2023; Tay et al., 2022; Wang et al., 2022; Zeng et al., 2024a, b; Zhuang et al., 2022). Docids could be predefined and remain static during training, making their careful design crucial for optimal retrieval performance (Tang et al., 2024b). We classify docids into two categories based on their generation methodology and abstraction level. The first category, referred to as content-derived docids, includes identifiers that are extracted directly from document elements such as titles (De Cao et al., 2020; Chen et al., 2022a, b; Lee et al., 2023), n-grams (Bevilacqua et al., 2022; Chen et al., 2023b; Li et al., 2023b, c; Wang et al., 2023), URLs (Ren et al., 2023; Zhou et al., 2023, 2022b; Ziems et al., 2023), and key terms (Zhang et al., 2023). These docids preserve surface-level textual characteristics and are closely tied to the original document content. In contrast, the second category, termed computationally-generated are derived using techniques like quantization (Chen et al., 2023a; Rajput et al., 2023; Zeng et al., 2024b; Zhou et al., 2022b; Zeng et al., 2024a) or hierarchical clustering algorithms (Mehta et al., 2023; Sun et al., 2023; Tay et al., 2022; Wang et al., 2022) to encode deeper semantics by abstracting raw document content into conceptual features.
@@
Our approach has two key phases. (i) First, we employ supervised fine-tuning (SFT) to train a language model (LM) capable of generating docids that are most relevant to a given query. (ii) Next, we directly refine the model through pairwise ranking, where the model learns to differentiate between relevant and irrelevant docids for a specific query based on labeled data. This refinement ensures that the retrieval system ranks documents based on their relevance, aligning the model more closely with the query. The SFT phase serves as a crucial pretraining step, aligning the model with initial relevance signals from training data and providing a strong foundation for the pairwise ranker to further fine-tune document ranking effectiveness.
@@
To promote reproducibility in GenIR, we open-source our codebase and checkpoints publicly available. The shared resources can be accessed at this GitHub repository
@@ 2. Related Work
Similarly, LTRGR (Li et al., 2023b) incorporates a learning-to-rank (LTR) framework to address the gap between docid generation and document ranking. It introduces an additional training phase where the model is optimized using a margin-based ranking loss, eliminating the need for a separate ranking step during inference. However, LTRGR focuses on optimizing passage ranking during the second phase, treating docid generation as a step toward this goal rather than fully integrating document-level relevance into the generative process. Consequently, the challenge of embedding document-level relevance directly into docid generation remains unaddressed. In contrast, DDRO integrates pairwise ranking directly into the generative model’s optimization pipeline, ensuring docid generation inherently aligns with document-level relevance.
@@ 3.2. Learning to Rank (L2R)
L2R aims at training models to rank documents based on their relevance to a given query (Burges, 2010; Cao et al., 2007; Wang et al., 2018). L2R methods can be classified into point-wise, pair-wise, and list-wise approaches based on their learning objectives. (i) Point-wise methods (Ibrahim and Carman, 2016) frame ranking as a classification problem by scoring individual documents independently: $L_{\text{point}}=\sum_{i}\mathcal{L}(\hat{s}(d_{i},q),s(d_{i},q))$, where $\hat{s}(d_{i},q)$ and $s(d_{i},q)$ denote the predicted relevance score and ground truth relevance score, respectively. In GenIR, generated probabilities serve as relevance scores, aligning with this approach (Tay et al., 2022). And the retrieval term in Eq. 2 belongs to this type. (ii) Pair-wise approaches  (Wang et al., 2018; Dammak et al., 2017; Burges, 2010; Cao et al., 2007) compare document pairs to determine relative preferences: $L_{\text{pair}}=\sum_{(d_{i},d_{j})}\log\left(1+\exp\left(-\left(\hat{s}(d_{i},q)-\hat{s}(d_{j},q)\right)\right)\right)$, where $d_{i}$ and $d_{j}$ are used as pairs to compare. Our DDRO follows this pair-wise framework using positive and negative docid pairs for learning relevance. (iii) List-wise approaches  (Xia et al., 2008; Lan et al., 2014), optimize the entire ranked list: $L_{\text{list}}=\sum_{q}\mathcal{L}\left(\text{Softmax}(\hat{s}(q,\hat{\pi})),\text{Softmax}(s(q,\pi))\right)$, where $\hat{\pi}$ and $\pi$ are the predicted and ground-truth lists, respectively. In GenIR, Tang et al. (2024a) introduces a position-aware list-level objective to learn the relevance. As we focus on pair-wise approaches, comparisons with list-wise methods are left for future work.
A fundamental challenge in GenIR stems from the inherent misalignment between the optimization objectives of autoregressive models and the overarching objectives of document ranking tasks. Training GenIR models solely to generate docids can be treat as the point-wise approach, which is often insufficient for achieving effective ranking. Addressing this challenge necessitates the development of a robust framework that enables GenIR models to directly learn to rank (Zhou et al., 2023; Li et al., 2023b; Zhou et al., 2024).
@@ 3.3. Reinforcement Learning from Relevance Feedback (RLRF)
GenRRL (Zhou et al., 2023) trains a reward model using relevance-annotated data derived from BM25 (Robertson and Zaragoza, 2009), DPR (Karpukhin et al., 2020b), and LLaMA-13b (Touvron et al., 2023). This reward model guides reinforcement learning to optimize the language model’s policy for generating high-reward outputs, with a KL divergence constraint ensuring alignment with the original supervised fine-tuned model or the reference policy $\pi^{\text{ref}}$. The optimization process involves supervised fine-tuning using negative log-likelihood, pairwise ranking loss for the reward model, and reinforcement learning techniques incorporating pointwise, pairwise, and listwise approaches to enhance ranking performance. While effective, this approach introduces considerable complexity, requiring the training of multiple models and sampling from the policy during training, which substantially increases computational costs.
Inspired by the work of  Rafailov et al. (2023), we propose direct document relevance optimization (DDRO), a streamlined approach designed to enhance the ranking capabilities of GenIR models. DDRO directly optimizes GenIR models to learn ranking without the need for explicit reward modeling or reinforcement learning. By simplifying the optimization process, DDRO provides an efficient alternative to reinforcement-based approaches, offering a more straightforward framework for improving the performance of GenIR models.
@@ 4.1. Docid Construction
URL and Title (TU). Titles in web search results are typically crafted to be descriptive, closely aligning with user search intent, while URLs often contain structured tokens, such as keywords or domains, that are highly indicative of relevance for web-based queries (Zhou et al., 2022b). The structure of the ULR reverses to prioritize semantically meaningful segments. When URLs lack sufficient descriptive content, a combination of the document’s title and domain name is used as an alternative identifier. Formally, this type of docid is defined as:
@@ 4.3. Direct L2R Optimization Using Relevance Feedback
where $\pi_{\theta}(docid|q)$ is the policy that is being optimized, while $\pi^{\text{ref}}(docid|q)$ is the reference policy, typically the fine-tuned model (SFT). This formulation ensures that the optimized model remains close to the reference policy while improving relevance-based ranking. The DDRO update efficiently guides the model toward producing outputs better aligned with relevance by utilizing pairwise comparisons, offering a simplified alternative to RL-based approaches. The gradient with respect to the model parameters $\theta$ is defined as in Eq. 7, where
@@ 5.1. Datasets and Evaluation Metrics
Following prior work (Zhou et al., 2023, 2022a, 2022b; Wang et al., 2023; Zhou et al., 2024), we assess model performance using standard document retrieval metrics: Recall (R@1/5/10) and Mean Reciprocal Rank (MRR@10). Statistical significance is determined using paired t-tests with a threshold of $p<0.05$.
@@ 5.2. Baselines
We evaluate our approach against three categories of baselines: term-based retrieval, dense retrieval, and generative retrieval.
@@
(i) DPR(Karpukhin et al., 2020b), which utilizes a BERT-based dual encoder to produce dense embeddings for queries and documents. PseudoQ (Tang et al., 2021) improves DPR by generating pseudo-queries using K-means clustering over document embeddings. (ii) ANCE(Xiong et al., 2020), a RoBERTa-based dual encoder that incorporates hard negatives retrieved from an asynchronously updated approximate nearest neighbor (ANN) index. (iii) RepBERT(Zhan et al., 2020), a BERT-based model that generates fixed-length contextualized embeddings, with query-document relevance computed via inner product similarity. (iv) Sentence-T5(Ni et al., 2022), which applies a T5-based architecture to generate sentence embeddings using encoder-only and encoder-decoder models with contrastive learning.
@@
(i) DSI(Tay et al., 2022), which represents docids using hierarchical k-means cluster IDs and trains with the DSI-Num objective. (ii) DSI-QG(Zhuang et al., 2022), which augments training data with synthetic queries generated using a query generation model (Nogueira and Lin, 2019) and represents documents with arbitrary unique numerical docids. (iii) NCI(Wang et al., 2022), which assigns semantically structured numeric docids paired with pseudo-queries. (iv) SEAL(De Cao et al., 2020), which retrieves docids represented as arbitrary n-grams extracted from document text using an FM-index. (v) Ultron(Zhou et al., 2022b), which employs keyword and semantic-based docids, using a three-stage training approach: general pre-training, search-oriented pre-training, and supervised fine-tuning. (vi) ROGER(Zhou et al., 2024), which transfers document relevance knowledge from a dense retriever to a generative retriever via knowledge distillation. (vii) MINDER(Li et al., 2023c), which assigns multiple identifiers, including titles, n-grams, and synthetic queries, to documents and pairs them for indexing. (viii) LTRGR(Li et al., 2023b), which trains on pairwise relevance objectives using margin-based ranking loss for optimization. (ix) GenRRL(Zhou et al., 2023), which incorporates pointwise, pairwise, and listwise relevance optimization through reinforcement learning, using document summaries and URLs as docids.
In this work, we exclude document summaries as docids due to their dependence on external summarization models, such as LLaMA-13b used in GenRRL. These models introduce preprocessing overhead and variability in identifier quality. Instead, DDRO employs product quantization (PQ) to generate compact, structured docids, ensuring consistency and scalability for large datasets.
@@ 6. Experimental Evaluation and Results
This section provides an empirical evaluation of the proposed DDRO framework, focusing on the following research questions:
@@ 6.1. Comparison with Reinforcement Learning-Based Methods
We can observe the following: (i) On MS MARCO, DDRO (TU) achieves the highest scores in early precision-focused metrics. Specifically, it outperforms GenRRL(Sum) by 15.06% in R@1 and 7.4% in MRR@10, highlighting its effectiveness in ranking the most relevant document at the top. While GenRRL (Sum) performs better in the R@10 metric, DDRO (TU) achieves comparable results with a simplified optimization process, avoiding the need for a reward model and reinforcement learning. (ii) On NQ, DDRO (PQ) outperforms GenRRL (Sum) by 34.69% in R@1 and 19.87% in MRR@10, further validating its effectiveness in retrieving relevant documents within the top ranks. GenRRL variants achieve higher R@10 scores, likely benefiting from multi-signal learning and listwise optimization strategies, potentially improving broader document retrieval. Furthermore, summary-based docids may aid performance on knowledge-intensive queries. (iii) DDRO (TU) and DDRO (PQ) show different trends across datasets: TU performs better on MS300K, while PQ excels on NQ. This disparity likely stems from higher document quality in NQ, where the rich semantic content allows the generated PQs to convey more meaningful information. In contrast, MS300K, which is derived from web search logs, contains noisy content such as ads, resulting in lower-quality PQs. Consequently, TU, which prioritizes keyword-based features, effectively captures the core content of MS300K, where the shorter, keyword-focused queries align well with surface-level signals such as titles and URLs.

### v2

@@ Abstract.
submissionid: 821copyright: rightsretainedjournalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730023isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Retrieval models and rankingccs: Information systems Learning to rankccs: Information systems Language models
@@ 1. Introduction
Building on the success of transformer-based pre-trained language models, recent research has explored various neural retrieval approaches (Yates et al., 2021): learned sparse retrieval (Formal et al., 2021a, b), dense retrieval (Karpukhin et al., 2020; Xiong et al., 2020; Zeng et al., 2022; Zhan et al., 2021), and cross-encoders (Nogueira and Cho, 2019). A new paradigm has recently been added to this palette, generative information retrieval (GenIR(Tay et al., 2022; Metzler et al., 2021). This approach uses pre-trained encoder-decoder models as differentiable search indexes (DSI). It has inspired the development of several models (Wang et al., 2022; Zhuang et al., 2022; Mehta et al., 2023; Bevilacqua et al., 2022; Zhou et al., 2022b; Ren et al., 2023; Zeng et al., 2024b, a; Tang et al., 2023b, 2024b, 2024a, 2025, 2024c)
@@
GenIR models represent documents as sequences of unique document identifiers (docids), generated autoregressively, where each token is conditioned on the query encoding and previously generated tokens. The generation process is controlled through (constrained) beam search (Mehta et al., 2023; Tay et al., 2022; Wang et al., 2022; Zeng et al., 2024a, b; Zhuang et al., 2022). Docids could be predefined and remain static during training, making their careful design crucial for optimal retrieval performance (Tang et al., 2024f). We classify docids into two categories based on their generation methodology and abstraction level. The first category, referred to as content-derived docids, includes identifiers that are extracted directly from document elements such as titles (De Cao et al., 2020; Chen et al., 2022a, b; Lee et al., 2023; Tang et al., 2024d, g, e, 2023a), n-grams (Bevilacqua et al., 2022; Chen et al., 2023b; Li et al., 2024, 2023b; Wang et al., 2023), URLs (Ren et al., 2023; Zhou et al., 2023, 2022b; Ziems et al., 2023), and key terms (Zhang et al., 2023). These docids preserve surface-level textual characteristics and are closely tied to the original document content. In contrast, the second category, termed computationally-generated are derived using techniques like quantization (Chen et al., 2023a; Rajput et al., 2023; Zeng et al., 2024b; Zhou et al., 2022b; Zeng et al., 2024a) or hierarchical clustering algorithms (Mehta et al., 2023; Sun et al., 2023; Tay et al., 2022; Wang et al., 2022) to encode deeper semantics by abstracting raw document content into conceptual features.
@@
Our approach has two key phases. (i) First, we employ supervised fine-tuning (SFT) to train a language model (LM) capable of generating docids that are most relevant to a given query. (ii) Next, we directly refine the model through pairwise ranking, where the model learns to differentiate between relevant and irrelevant docids for a specific query based on labeled data. This refinement ensures that the retrieval system ranks documents based on their relevance, aligning the model more closely with the query. The SFT phase serves as a pretraining step, aligning the model with initial relevance signals from training data and providing a strong foundation for the pairwise ranker to further fine-tune document ranking effectiveness.
@@
To promote reproducibility in GenIR, we open-source our codebase and make checkpoints publicly available. [^1] https://github.com/kidist-amde/DDRO-Direct-Document-Relevance-Optimization/tree/main
@@ 2. Related Work
Similarly, LTRGR (Li et al., 2024) incorporates a learning-to-rank (LTR) framework to address the gap between docid generation and document ranking. It introduces an additional training phase where the model is optimized using a margin-based ranking loss, eliminating the need for a separate ranking step during inference. However, LTRGR focuses on optimizing passage ranking during the second phase, treating docid generation as a step toward this goal rather than fully integrating document-level relevance into the generative process. Consequently, the challenge of embedding document-level relevance directly into docid generation remains unaddressed. In contrast, DDRO integrates pairwise ranking directly into the generative model’s optimization pipeline, ensuring docid generation inherently aligns with document-level relevance.
@@ 3.2. Learning to Rank (L2R)
L2R aims at training models to rank documents based on their relevance to a given query (Burges, 2010; Cao et al., 2007; Wang et al., 2018). L2R methods can be classified into point-wise, pair-wise, and list-wise approaches based on their learning objectives. (i) Point-wise methods (Ibrahim and Carman, 2016) frame ranking as a classification problem by scoring individual documents independently: $L_{\text{point}}=\sum_{i}\mathcal{L}(\hat{s}(d_{i},q),s(d_{i},q))$, where $\hat{s}(d_{i},q)$ and $s(d_{i},q)$ denote the predicted relevance score and ground truth relevance score, respectively. In GenIR, generated probabilities serve as relevance scores, aligning with this approach (Tay et al., 2022). And the retrieval term in Eq. 2 belongs to this type. (ii) Pair-wise approaches  (Wang et al., 2018; Dammak et al., 2017; Burges, 2010; Cao et al., 2007) compare document pairs to determine relative preferences: $L_{\text{pair}}=\sum_{(d_{i},d_{j})}\log\left(1+\exp\left(-\left(\hat{s}(d_{i},q)-\hat{s}(d_{j},q)\right)\right)\right)$, where $d_{i}$ and $d_{j}$ are used as pairs to compare. DDRO shares similarities with traditional pairwise L2R methods such as RankNet (Burges et al., 2005) and LambdaRank (Burges et al., 2006), in that it optimizes a margin between relevant and non-relevant documents. It differs in that the ranking signal is used to supervise the generation of structured docid sequences via a generative decoder. Unlike typical L2R approaches that score documents retrieved by an external system, DDRO learns to produce docids directly, making it end-to-end generative. This integration of sequence modeling and pairwise supervision is a key distinction from prior L2R pipelines. (iii) List-wise approaches  (Xia et al., 2008; Lan et al., 2014), optimize the entire ranked list: $L_{\text{list}}=\sum_{q}\mathcal{L}\left(\text{Softmax}(\hat{s}(q,\hat{\pi})),\text{Softmax}(s(q,\pi))\right)$, where $\hat{\pi}$ and $\pi$ are the predicted and ground-truth lists, respectively. In GenIR, Tang et al. (2024b) introduces a position-aware list-level objective to learn the relevance. As we focus on pair-wise approaches, comparisons with list-wise methods are left for future work.
A fundamental challenge in GenIR stems from the inherent misalignment between the optimization objectives of autoregressive models and the overarching objectives of document ranking tasks. Training GenIR models solely to generate docids can be treat as the point-wise approach, which is often insufficient for achieving effective ranking. Addressing this challenge necessitates the development of a robust framework that enables GenIR models to directly learn to rank (Zhou et al., 2023; Li et al., 2024; Zhou et al., 2024).
@@ 3.3. Reinforcement Learning from Relevance Feedback (RLRF)
GenRRL (Zhou et al., 2023) trains a reward model using relevance-annotated data derived from BM25 (Robertson and Zaragoza, 2009), DPR (Karpukhin et al., 2020), and LLaMA-13b (Touvron et al., 2023). This reward model guides reinforcement learning to optimize the language model’s policy for generating high-reward outputs, with a KL divergence constraint ensuring alignment with the original supervised fine-tuned model or the reference policy $\pi^{\text{ref}}$. The optimization process involves supervised fine-tuning using negative log-likelihood, pairwise ranking loss for the reward model, and reinforcement learning techniques incorporating pointwise, pairwise, and listwise approaches to enhance ranking performance. While effective, this approach introduces considerable complexity, requiring the training of multiple models and sampling from the policy during training, which substantially increases computational costs.
Inspired by the work of Rafailov et al. (2023), we propose direct document relevance optimization (DDRO), a streamlined approach designed to enhance the ranking capabilities of GenIR models. DDRO directly optimizes GenIR models to learn document-level ranking without relying on explicit reward modeling or reinforcement learning. While our optimization is inspired by the DPO framework (Rafailov et al., 2023), its adaptation to GenIR is non-trivial. Unlike preference alignment for open-ended generation, our task involves optimizing structured docid generation under beam decoding constraints. Additionally, our method differs in both architecture (encoder-decoder vs. decoder-only) and objective (document ranking vs. preference alignment), requiring novel integration into GenIR pipelines. To the best of our knowledge, DDRO is the first method to apply preference-style optimization directly to generative document retrieval by extending DPO-style training to constrained generative settings, leveraging pairwise query-docid relevance supervision and constrained decoding.
@@ 4.1. Docid Construction
URL and Title (TU). Titles in web search results are typically crafted to be descriptive, closely aligning with user search intent, while URLs often contain structured tokens, such as keywords or domains, that are highly indicative of relevance for web-based queries (Zhou et al., 2022b). The structure of the URL reverses to prioritize semantically meaningful segments. When a URL lacks descriptive content (e.g., uses numeric IDs or generic paths), we fall back to a combination of the document’s title and domain name as an alternative identifier. Formally, this docid variant is defined as:
@@ 4.3. Direct L2R Optimization Using Relevance Feedback
where $\pi_{\theta}(docid|q)$ is the policy that is being optimized, while $\pi^{\text{ref}}(docid|q)$ is the reference policy, typically the fine-tuned model (SFT). This formulation ensures that the optimized model remains close to the reference policy while improving relevance-based ranking. The DDRO update guides the model toward producing outputs better aligned with relevance by utilizing pairwise comparisons, offering a simplified alternative to RL-based approaches. The gradient w.r.t. the model parameters $\theta$ is defined as in Eq. 7, where
@@ 5.1. Datasets and Evaluation Metrics
Following (Zhou et al., 2023, 2022a, 2022b; Wang et al., 2023; Zhou et al., 2024), we assess model performance using standard document retrieval metrics: Recall (R@1/5/ 10) and Mean Reciprocal Rank (MRR@10). Statistical significance is determined using paired t-tests with a threshold of $p<0.05$.
@@ 5.2. Baselines
We evaluate our approach against three types of baseline: term-based retrieval, dense retrieval, and generative retrieval.
@@
(i) DPR(Karpukhin et al., 2020), which utilizes a BERT-based dual encoder to produce dense embeddings for queries and documents. PseudoQ (Tang et al., 2021) improves DPR by generating pseudo-queries using K-means clustering over document embeddings. (ii) ANCE(Xiong et al., 2020), a RoBERTa-based dual encoder that incorporates hard negatives retrieved from an asynchronously updated approximate nearest neighbor (ANN) index. (iii) RepBERT(Zhan et al., 2020), a BERT-based model that generates fixed-length contextualized embeddings, with query-document relevance computed via inner product similarity. (iv) Sentence-T5(Ni et al., 2022), which applies a T5-based architecture to generate sentence embeddings using encoder-only and encoder-decoder models with contrastive learning.
@@
(i) DSI(Tay et al., 2022), which represents docids using hierarchical k-means cluster IDs and trains with the DSI-Num objective. (ii) DSI-QG(Zhuang et al., 2022), which augments training data with synthetic queries generated using a query generation model (Nogueira and Lin, 2019) and represents documents with arbitrary unique numerical docids. (iii) NCI(Wang et al., 2022), which assigns semantically structured numeric docids paired with pseudo-queries. (iv) SEAL(De Cao et al., 2020), which retrieves docids represented as arbitrary n-grams extracted from document text using an FM-index. (v) Ultron(Zhou et al., 2022b), which employs keyword and semantic-based docids, using a three-stage training approach: general pre-training, search-oriented pre-training, and supervised fine-tuning. (vi) ROGER(Zhou et al., 2024), which transfers document relevance knowledge from a dense retriever to a generative retriever via knowledge distillation. (vii) MINDER(Li et al., 2023b), which assigns multiple identifiers, including titles, n-grams, and synthetic queries, to documents and pairs them for indexing. (viii) LTRGR(Li et al., 2024), which trains on pairwise relevance objectives using margin-based ranking loss for optimization. (ix) GenRRL(Zhou et al., 2023), which incorporates pointwise, pairwise, and listwise relevance optimization through reinforcement learning, using document summaries and URLs as docids.
We exclude document summaries as docids due to their dependence on external summarization models, such as LLaMA-13b used in GenRRL. These models introduce preprocessing overhead and variability in identifier quality. Instead, DDRO employs product quantization (PQ) to generate compact, structured docids, ensuring consistency and scalability.
@@ 6. Experimental Evaluation and Results
Our evaluation of DDRO focuses on the following questions:
@@ 6.1. Comparison with Reinforcement Learning-Based Methods
We observe the following: (i) On MS MARCO, DDRO (TU) achieves the highest scores in early precision-focused metrics. Specifically, it outperforms GenRRL(Sum) by 15.06% in R@1 and 7.4% in MRR@10, highlighting its effectiveness in ranking the most relevant document at the top. While GenRRL (Sum) performs better in the R@10 metric, DDRO (TU) achieves comparable results with a simplified optimization process, avoiding the need for a reward model and reinforcement learning. (ii) On NQ, DDRO (PQ) outperforms GenRRL (Sum) by 34.69% in R@1 and 19.87% in MRR@10, further validating its effectiveness in retrieving relevant documents within the top ranks. GenRRL variants achieve higher R@10 scores, likely benefiting from multi-signal learning and listwise optimization strategies, potentially improving broader document retrieval. Summary-based docids may aid performance on knowledge-intensive queries. (iii) DDRO (TU) and DDRO (PQ) show different trends across datasets: TU performs better on MS300K, while PQ excels on NQ. This disparity likely stems from higher document quality in NQ, where the rich semantic content allows the generated PQs to convey more meaningful information. In contrast, MS300K, which is derived from web search logs, contains noisy content such as ads, resulting in lower-quality PQs. Consequently, TU, which prioritizes keyword-based features, effectively captures the core content of MS300K, where the shorter, keyword-focused queries align well with surface-level signals such as titles and URLs.